# Test time training enhances in-context learning of nonlinear functions

### v1

@@ 4 Synthetic experiment
We conducted a numerical experiment to examine the effectiveness of TTT compared to ICL. To evaluate whether our theory holds with real-world settings, we trained a GPT-2 model (Radford et al., 2019) to learn the Gaussian single-index functions. See Appendix I for further experimental details.
The dimensions were set to $d=r=16$. For each task $t$, the data was generated as follows: ${\bm{x}}_{1}^{t},\cdots,{\bm{x}}_{N}^{t},{\bm{x}}^{t}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,{\bm{I}}_{d})$, $\beta^{t}\sim\mathrm{Unif}(\mathbb{S}^{d-1})$ and $y_{i}^{t}=\sigma_{*}^{t}(\left\langle\beta,{\bm{x}}_{i}\right\rangle)=\frac{1}{\sqrt{3!}}\mathrm{He}_{3}(\left\langle\beta,{\bm{x}}_{i}\right\rangle)+\frac{c^{t}}{\sqrt{4!}}\mathrm{He}_{4}(\left\langle\beta,{\bm{x}}_{i}\right\rangle)$, where $c^{t}\sim\mathrm{Unif}(-0.5,0.5)$. We compared the following two settings:
@@
    In-context learning, configurated as Garg et al. (2023), Oko et al. (2024) and Nishikawa et al. (2025). We first pretrained the model with the data $({\bm{X}}^{t},{\bm{y}}^{t},{\bm{x}}^{t},y^{t})_{t=1}^{T_{pt}}$, then we calculated the predictive loss.
@@
    Test-time training: We adapted the same pre-trained model as we used in the evaluation of ICL by fine-tuning the LoRA parameters applied to the attention and MLP layers.
Figure  1(a) highlights the result. Contrary to our initial expectation, the pretrained model’s ICL ability is powerful enough to adapt to the varying link function, while TTT does not lead to significant improvement of the prediction.
This result spurred us to analyze the limitation of ICL and the true potential of TTT in more challenging tasks: we changed the distribution of link function as $y_{i}^{t}=\frac{c_{3}^{t}}{\sqrt{3!}}\mathrm{He}_{3}(\left\langle\beta,{\bm{x}}_{i}\right\rangle)+\frac{c_{4}^{t}}{\sqrt{4!}}\mathrm{He}_{4}(\left\langle\beta,{\bm{x}}_{i}\right\rangle)$, where $c_{3}^{t}\sim\mathrm{Unif}(0.5,1.5),c_{4}^{t}\sim\mathrm{Unif}(-0.5,0.5)$, only in test-time. The result, summarized in Figure 1(b), clearly shows the advantage of TTT over ICL. As the data length grows, the accuracy of TTT steadily improves and surpasses that of ICL. This highlights an important benefit of TTT: TTT can adapt to distribution shifts, while ICL fails to do so.
(a) In-distribution setting: The distribution of the link function is consistent between pretraining and test-time.
(b) Out-of-distribution setting: The distribution of the link function at test-time differs from that of pre-training
Figure 1: The predictive error of In-context learning (ICL) and Test-time training (TTT) for a pretrained GPT-2 model on single-index polynomials (see Section 4 for details). While (a) shows that ICL effectively adapts to varying link functions for in-distribution tasks, (b) demonstrates that TTT outperforms ICL under distribution shift, showing its superior adaptability.
@@ 5 Conclusion
  • The result of the numerical experiment demonstrates that the ICL model’s loss decreases as we increase $n$ in in-distribution cases. This finding is contrary to our theoretical expectation that ICL cannot adapt to the change in link function. Future work could clarify whether ICL without TTT can manage varying link functions.
@@
Ackonwledgement
@@ Appendix C Gradient calculation
For the sake of simplicity, we write $\sigma_{*}^{\mathrm{test}}$ as $\sigma_{*}$. Let $\Gamma_{u}=\Gamma^{*}+uu^{\mathrm{T}}$, ${\bm{x}}_{i}^{*}=\sqrt{r}\Gamma^{*}{\bm{x}}_{i}$, $\beta^{*}=\sqrt{r}\Gamma^{*}\beta$ and $f_{\mathrm{IC}}({\bm{x}})=\sum_{j=1}^{m}a_{j}\sigma\left(v_{j}\frac{\sum_{i=1}^{N_{1}}y_{i}\mathrm{e}^{y_{i}/\rho}\mathrm{e}^{{\bm{x}}_{i}^{*\top}\Gamma_{u}{\bm{x}}/\rho}}{\sum_{i=1}^{N_{1}}\mathrm{e}^{y_{i}/\rho}\mathrm{e}^{{\bm{x}}_{i}^{*\top}\Gamma_{u}{\bm{x}}/\rho}}+b_{j}\right)$. Moreover, let us define $A_{i},B_{i}$ as $\bar{\sigma}_{*}(z)\exp((\bar{\sigma}_{*}(z))=\sum_{i\geq 0}\frac{A_{i}}{i!}z^{i}$ and $\exp((\bar{\sigma}_{*}(z))=\sum_{i\geq 0}\frac{B_{i}}{i!}z^{i}$, where
@@
Lemma 9.
Take $\rho=\Theta((\log{d})^{C_{\rho}})$ where $C_{\rho}$ is a constant sufficiently large. Suppose that $\bm{u}$} satisfies $\|u\|=C_{u}\leq 1$ and $\left\langle{\bm{u}},{\bm{x}}_{i}^{*}\right\rangle=C_{u}\tilde{O}_{d}(1)$. Then
@@
Note that from Lemma 29, the condition $\left\langle{\bm{u}},{\bm{x}}_{i}^{*}\right\rangle=C_{u}\tilde{O}_{d}(1)$ is satisfied at the initialization.
@@
First, from the assumption $\left\langle{\bm{u}},{\bm{x}}_{i}^{*}\right\rangle=C_{u}\tilde{O}_{d}(1)$ with $C_{u}\leq 1$ and Lemma 25, we can see that
@@
$\mathbb{E}_{{\bm{x}}_{1},\zeta_{1}}[\mathrm{e}^{y_{1}/\rho}\mathrm{e}^{{\bm{x}}_{1}^{*\top}\Gamma_{u}{\bm{x}}/\rho}{\bm{x}}_{1}^{*}]=$ $k_{\rho}\mathbb{E}_{{\bm{x}}_{1}}[\mathrm{e}^{\sigma_{*}(\left\langle\beta,{\bm{x}}_{1}\right\rangle)}\mathrm{e}^{{\bm{x}}_{1}^{*\top}\Gamma_{u}{\bm{x}}/\rho}{\bm{x}}_{1}^{*}]$ (1)
@@
$\mathbb{E}_{{\bm{x}}_{1},\zeta_{1}}[\mathrm{e}^{y_{1}/\rho}\mathrm{e}^{{\bm{x}}_{1}^{*\top}\Gamma_{u}{\bm{x}}/\rho}{\bm{x}}_{1}^{*}]$
@@
$\mathbb{E}_{{\bm{x}}_{1},\zeta_{1}}\left[\frac{y_{1}}{\rho}\mathrm{e}^{y_{1}/\rho}\mathrm{e}^{{\bm{x}}_{1}^{*\top}\Gamma_{u}{\bm{x}}/\rho}{\bm{x}}_{1}^{*}\right]$
@@
$\mathbb{E}_{{\bm{x}}_{1},\zeta_{1}}\left[\mathrm{e}^{y_{1}/\rho}\mathrm{e}^{{\bm{x}}_{1}^{*\top}\Gamma_{u}{\bm{x}}/\rho}\right]=k_{\rho}\exp\left(\frac{\|\sqrt{r}\Gamma^{*\top}\Gamma_{u}{\bm{x}}\|^{2}}{2\rho^{2}}\right)\left(\sum_{i\geq 0}\frac{B_{i}}{i!}z^{i}\right),$
@@
$\mathbb{E}_{{\bm{x}}_{1},\zeta_{1}}\left[\frac{y_{1}}{\rho}\mathrm{e}^{y_{1}/\rho}\mathrm{e}^{{\bm{x}}_{1}^{*\top}\Gamma_{u}{\bm{x}}/\rho}\right]=k_{\rho}\exp\left(\frac{\|\sqrt{r}\Gamma^{*\top}\Gamma_{u}{\bm{x}}\|^{2}}{2\rho^{2}}\right)\left(\sum_{i\geq 0}\frac{A_{i}}{i!}z^{i}\right)+k_{\rho}^{\prime}\exp\left(\frac{\|\sqrt{r}\Gamma^{*\top}\Gamma_{u}{\bm{x}}\|^{2}}{2\rho^{2}}\right)\left(\sum_{i\geq 0}\frac{B_{i}}{i!}z^{i}\right).$
@@
$\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\frac{y_{i}}{\rho}\mathrm{e}^{y_{i}/\rho}\mathrm{e}^{{\bm{x}}_{i}^{*\top}\Gamma_{u}{\bm{x}}/\rho}=\mathbb{E}_{{\bm{x}}_{1},\zeta_{1}}\left[\frac{y_{1}}{\rho}\mathrm{e}^{y_{1}/\rho}\mathrm{e}^{{\bm{x}}_{1}^{*\top}\Gamma_{u}{\bm{x}}/\rho}\right]+\tilde{O}(N_{1}^{-1/2}),$
@@
$\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\mathrm{e}^{y_{i}/\rho}\mathrm{e}^{{\bm{x}}_{i}^{*\top}\Gamma_{u}{\bm{x}}/\rho}=\mathbb{E}_{{\bm{x}}_{1},\zeta_{1}}\left[\mathrm{e}^{y_{1}/\rho}\mathrm{e}^{{\bm{x}}_{1}^{*\top}\Gamma_{u}{\bm{x}}/\rho}\right]+\tilde{O}(N_{1}^{-1/2}).$
Moreover, noting that $\sqrt{r}\Gamma^{*}{\bm{x}}_{i}^{*}=\tilde{O}(\sqrt{r})$ (see Lemma 28), from Lemma 30, we have
$\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\sqrt{r}\Gamma^{*}\frac{y_{i}}{\rho}\mathrm{e}^{y_{i}/\rho}\mathrm{e}^{{\bm{x}}_{i}^{*\top}\Gamma_{u}{\bm{x}}/\rho}{\bm{x}}_{i}^{*}=\sqrt{r}\Gamma^{*}\mathbb{E}_{{\bm{x}}_{1},\zeta_{1}}\left[\frac{y_{1}}{\rho}\mathrm{e}^{y_{1}/\rho}\mathrm{e}^{{\bm{x}}_{1}^{*\top}\Gamma_{u}{\bm{x}}/\rho}{\bm{x}}_{1}^{*}\right]+\tilde{O}(r^{1/2}N_{1}^{-1/2}),$
@@
$\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\sqrt{r}\Gamma^{*}\mathrm{e}^{y_{i}/\rho}\mathrm{e}^{{\bm{x}}_{i}^{*\top}\Gamma_{u}{\bm{x}}/\rho}{\bm{x}}_{i}^{*}=\sqrt{r}\Gamma^{*}\mathbb{E}_{{\bm{x}}_{1},\zeta_{1}}\left[\mathrm{e}^{y_{1}/\rho}\mathrm{e}^{{\bm{x}}_{1}^{*\top}\Gamma_{u}{\bm{x}}/\rho}{\bm{x}}_{1}^{*}\right]+\tilde{O}(r^{1/2}N_{1}^{-1/2}).$
@@
$\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\frac{y_{i}}{\rho}\mathrm{e}^{y_{i}/\rho}\mathrm{e}^{{\bm{x}}_{i}^{*\top}\Gamma_{u}{\bm{x}}/\rho}\left\langle{\bm{x}}_{i}^{*},{\bm{u}}\right\rangle=\mathbb{E}_{{\bm{x}}_{1},\zeta_{1}}\left[\frac{y_{1}}{\rho}\mathrm{e}^{y_{1}/\rho}\mathrm{e}^{{\bm{x}}_{1}^{*\top}\Gamma_{u}{\bm{x}}/\rho}\left\langle{\bm{x}}_{1}^{*},{\bm{u}}\right\rangle\right]+\tilde{O}(C_{u}N_{1}^{-1/2}),$
@@
$\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\mathrm{e}^{y_{i}/\rho}\mathrm{e}^{{\bm{x}}_{i}^{*\top}\Gamma_{u}{\bm{x}}/\rho}\left\langle{\bm{x}}_{i}^{*},{\bm{u}}\right\rangle=\mathbb{E}_{{\bm{x}}_{1},\zeta_{1}}\left[\mathrm{e}^{y_{1}/\rho}\mathrm{e}^{{\bm{x}}_{1}^{*\top}\Gamma_{u}{\bm{x}}/\rho}\left\langle{\bm{x}}_{1}^{*},{\bm{u}}\right\rangle\right]+\tilde{O}(C_{u}N_{1}^{-1/2}).$
@@
Moreover, from Lemma 9, it holds that
@@
where $\sqrt{r}\Gamma^{*}{\bm{x}}=\tilde{O}(\sqrt{r})$ (See Lemma 28) and $\gamma({\bm{x}},y)=\tilde{O}(1)$ with high probability, which means that $\sqrt{r}\Gamma^{*}\frac{\partial f_{\mathrm{IC}}}{\partial{\bm{u}}}=\tilde{O}_{p}(\alpha mC_{u}\sqrt{r})$. Therefore, we obtain
@@
Moreover, $\mathbb{E}_{{\bm{x}}}[\nu]=0$ by definition, and $\|\nu\|$ has sub-Weibull tail from Lemma 10. Thus, what remains is to calculate $\mathbb{E}_{{\bm{x}}}[y\cdot\sqrt{r}\Gamma^{*}\nabla_{\bm{u}}{f_{\mathrm{IC}}}({\bm{x}})]$:
@@
and from Lemma 31, it holds that
@@
with high probability. Finally, we evaluate the term $\mathbb{E}_{{\bm{x}}}[\alpha L_{m}\sigma_{*}(\left\langle\beta,{\bm{x}}\right\rangle)(\sqrt{r}\Gamma^{*}{\bm{x}}\cdot n_{1}+\left\langle{\bm{x}},{\bm{u}}\right\rangle{\bm{n}}_{2})]$. From Corollary 17 of Oko et al. (2024), we have $\sigma_{*}(\left\langle\beta,{\bm{x}}\right\rangle)=O_{p}(\log^{\deg(\sigma_{*})/2}{d})$. Moreover, from Lemma 9, it holds that $n_{1}=\tilde{O}(C_{u}\sqrt{\frac{1}{N_{1}}})$ and ${\bm{n}}_{2}=\tilde{O}(\sqrt{\frac{r}{N_{1}}})$ with high probability. Therefore, we arrive at

### v2

@@ 4 Synthetic experiment
We conducted a numerical experiment to examine the effectiveness of TTT compared to ICL. To align the experimental setup with our theoretical framework, we employed a 2-layer GPT-2 model (Radford et al., 2019) to learn Gaussian single-index functions. See Appendix I for further experimental details.
The ambient and intrinsic dimensions were set to $d=r=4$. For each task $t$, the data was generated as follows: ${\bm{x}}_{1}^{t},\cdots,{\bm{x}}_{N}^{t},{\bm{x}}^{t}\overset{\text{i.i.d.}}{\sim}\mathcal{N}(0,{\bm{I}}_{d})$, $\beta^{t}\sim\mathrm{Unif}(\mathbb{S}^{r-1})$ and $y_{i}^{t}=\sigma_{*}^{t}(\left\langle\beta,{\bm{x}}_{i}\right\rangle)=\frac{1}{\sqrt{3!}}\mathrm{He}_{3}(\left\langle\beta,{\bm{x}}_{i}\right\rangle)+\frac{c^{t}}{\sqrt{4!}}\mathrm{He}_{4}(\left\langle\beta,{\bm{x}}_{i}\right\rangle)$, where $c^{t}\sim\mathrm{Unif}(-0.5,0.5)$. We compared the following two settings:
@@
    In-context learning, configurated as Garg et al. (2023), Oko et al. (2024) and Nishikawa et al. (2025), where the model makes predictions solely based on the prompt without weight updates. The model was pretrained on a standard task where the link function $\sigma_{*}^{t}$ varied across all prompts.
@@
    Test-time training: We introduced a strategic pretraining phase: to facilitate the convergence of the attention mechanism on the target subspace, we fixed the link function as $\sigma_{*}^{t}=\frac{1}{\sqrt{3!}}\mathrm{He}_{3}$ during the pretraining. Once pretrained, the model was evaluated on the same challenging task as ICL, where the link function varies and requires test-time adaptation via LoRA.
Figure  1 highlights the result. The predictive error of vanilla ICL remains high and fails to decrease with increasing context length, suggesting that ICL alone cannot adapt to shifts in the link function. In contrast, TTT exhibits steady convergence, with the error diminishing significantly as more context examples are provided. While TTT shows initial instability at very short context lengths due to high-learning-rate updates, it ultimately demonstrates a superior ability to fit task-specific non-linearities. Overall, this result demonstrates a great advantage of TTT: TTT allows the model to adapt to task-specific link functions.
Figure 1: The predictive error of In-context learning (ICL) and Test-time training (TTT) for a pretrained GPT-2 model on single-index polynomials (see Section 4 for details). While vanilla ICL fails to adapt to task-specific link functions, the error for TTT steadily decreases as context length grows. This highlights TTT’s ability to fit the model’s nonlinearity at inference time.
To further assess the capabilities of TTT, we examined its ability to leverage low-dimensional task structures. We fixed the intrinsic dimension at $r=4$ and compared the predictive error across varying ambient dimensions $d\in\{4,8,16\}$. As illustrated in Figure 2, the model performance remains virtually unchanged despite the fourfold increase in the ambient dimension. This empirical result provides strong evidence for our main theorem: the sample complexity of TTT-equipped transformers is governed by the intrinsic dimension $r$ rather than the ambient dimension $d$, showcasing the model’s ability to exploit low-dimensional task structures effectively.
Figure 2: Predictive error of TTT across different ambient dimensions $d\in\{4,8,16\}$ with a fixed intrinsic dimension $r=4$. The overlap of the convergence curves demonstrates that TTT is robust to increases in the ambient dimension, empirically verifying that sample complexity scales with the intrinsic dimensionality $r$.
@@ 5 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Acknowledgement
@@ Appendix B Pretraining
$g(\Gamma^{*},{\bm{X}}_{N_{1}},{\bm{y}}_{N_{1}},{\bm{x}})=P_{1}^{\prime}+P_{2}^{\prime}\left(\frac{\left\langle{\bm{x}},\beta\right\rangle}{\sqrt{r}}\right)^{\mathrm{ge}(\sigma_{*})}+n_{3},$
@@ Appendix C Gradient calculation
For the sake of simplicity, we write $\sigma_{*}^{\mathrm{test}}$ as $\sigma_{*}$, and $f_{\mathrm{IC}}(\Gamma,{\bm{X}}_{N},{\bm{y}}_{N},{\bm{x}})=\sum_{j=1}^{m}a_{j}\sigma\left(v_{j}\frac{\sum_{i=1}^{N_{1}}y_{i}\mathrm{e}^{y_{i}/\rho}\mathrm{e}^{{\bm{x}}_{i}^{*\top}\Gamma_{u}{\bm{x}}/\rho}}{\sum_{i=1}^{N_{1}}\mathrm{e}^{y_{i}/\rho}\mathrm{e}^{{\bm{x}}_{i}^{*\top}\Gamma_{u}{\bm{x}}/\rho}}+b_{j}\right)$ as $f_{\mathrm{IC}}({\bm{x}})$. Let $\Gamma_{u}=\Gamma^{*}+uu^{\mathrm{T}}$, ${\bm{x}}_{i}^{*}=\sqrt{r}\Gamma^{*}{\bm{x}}_{i}$, $\beta^{*}=\sqrt{r}\Gamma^{*}\beta$. We define $L_{m}$ as $|\{j\in\{1,2,\dots,m\}\mid v_{j}(0)=1\}|$. Moreover, let us define $A_{i},B_{i}$ as $\bar{\sigma}_{*}(z)\exp((\bar{\sigma}_{*}(z))=\sum_{i\geq 0}\frac{A_{i}}{i!}z^{i}$, and $\exp((\bar{\sigma}_{*}(z))=\sum_{i\geq 0}\frac{B_{i}}{i!}z^{i}$, where
@@
Lemma C.1.
Take $\rho=\Theta((\log{d})^{C_{\rho}})$ where $C_{\rho}$ is a constant sufficiently large. Let $\alpha$ be the initial scale of the MLP parameters $a_{j}$. Suppose that $\bm{u}$} satisfies $\|u\|=C_{u}\leq 1$ and $\left\langle{\bm{u}},{\bm{x}}_{i}^{*}\right\rangle=C_{u}\tilde{O}_{d}(1)$. Then
@@
Note that from Lemma H.5, the condition $\left\langle{\bm{u}},{\bm{x}}_{i}^{*}\right\rangle=C_{u}\tilde{O}_{d}(1)$ is satisfied at the initialization.
@@
First, from the assumption $\left\langle{\bm{u}},{\bm{x}}_{i}^{*}\right\rangle=C_{u}\tilde{O}_{d}(1)$ with $C_{u}\leq 1$ and Lemma H.1, we can see that
@@
$\mathbb{E}_{{\bm{x}}_{1},\zeta_{1}}[\mathrm{e}^{\bar{\sigma}_{*}(\left\langle\beta,{\bm{x}}_{1}\right\rangle)/\rho+\zeta_{1}/\rho}\mathrm{e}^{{\bm{x}}_{1}^{*\top}\Gamma_{u}{\bm{x}}/\rho}{\bm{x}}_{1}^{*}]=$ $k_{\rho}\mathbb{E}_{{\bm{x}}_{1}}[\mathrm{e}^{\sigma_{*}(\left\langle\beta,{\bm{x}}_{1}\right\rangle)}\mathrm{e}^{{\bm{x}}_{1}^{*\top}\Gamma_{u}{\bm{x}}/\rho}{\bm{x}}_{1}^{*}]$ (1)
@@
$\mathbb{E}_{{\bm{x}}_{1},\zeta_{1}}[\mathrm{e}^{\bar{\sigma}_{*}(\left\langle{\bm{x}}_{1},\beta\right\rangle)/\rho+\zeta_{1}/\rho}\mathrm{e}^{{\bm{x}}_{1}^{*\top}\Gamma_{u}{\bm{x}}/\rho}{\bm{x}}_{1}^{*}]$
@@
$\mathbb{E}_{{\bm{x}}_{1},\zeta_{1}}\left[\frac{\bar{\sigma}_{*}(\left\langle{\bm{x}}_{1},\beta\right\rangle)+\zeta_{1}}{\rho}\mathrm{e}^{\bar{\sigma}_{*}(\left\langle{\bm{x}}_{1},\beta\right\rangle)/\rho+\zeta_{1}/\rho}\mathrm{e}^{{\bm{x}}_{1}^{*\top}\Gamma_{u}{\bm{x}}/\rho}{\bm{x}}_{1}^{*}\right]$
@@
$\mathbb{E}_{{\bm{x}}_{1},\zeta_{1}}\left[\mathrm{e}^{\bar{\sigma}_{*}(\left\langle{\bm{x}}_{1},\beta\right\rangle)/\rho+\zeta_{1}/\rho}\mathrm{e}^{{\bm{x}}_{1}^{*\top}\Gamma_{u}{\bm{x}}/\rho}\right]=k_{\rho}\exp\left(\frac{\|\sqrt{r}\Gamma^{*\top}\Gamma_{u}{\bm{x}}\|^{2}}{2\rho^{2}}\right)\left(\sum_{i\geq 0}\frac{B_{i}}{i!}z^{i}\right),$
@@
$\mathbb{E}_{{\bm{x}}_{1},\zeta_{1}}\left[\frac{\bar{\sigma}_{*}(\left\langle{\bm{x}}_{1},\beta\right\rangle)+\zeta_{1}}{\rho}\mathrm{e}^{\bar{\sigma}_{*}(\left\langle{\bm{x}}_{1},\beta\right\rangle)/\rho+\zeta_{1}/\rho}\mathrm{e}^{{\bm{x}}_{1}^{*\top}\Gamma_{u}{\bm{x}}/\rho}\right]$
$=$ $k_{\rho}\exp\left(\frac{\|\sqrt{r}\Gamma^{*\top}\Gamma_{u}{\bm{x}}\|^{2}}{2\rho^{2}}\right)\left(\sum_{i\geq 0}\frac{A_{i}}{i!}z^{i}\right)+k_{\rho}^{\prime}\exp\left(\frac{\|\sqrt{r}\Gamma^{*\top}\Gamma_{u}{\bm{x}}\|^{2}}{2\rho^{2}}\right)\left(\sum_{i\geq 0}\frac{B_{i}}{i!}z^{i}\right).$
@@
$\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\frac{y_{i}}{\rho}\mathrm{e}^{y_{i}/\rho}\mathrm{e}^{{\bm{x}}_{i}^{*\top}\Gamma_{u}{\bm{x}}/\rho}=\mathbb{E}_{{\bm{x}}_{1},\zeta_{1}}\left[\frac{\bar{\sigma}_{*}(\left\langle{\bm{x}}_{1},\beta\right\rangle)+\zeta_{1}}{\rho}\mathrm{e}^{\bar{\sigma}_{*}(\left\langle{\bm{x}}_{1},\beta\right\rangle)/\rho+\zeta_{1}/\rho}\mathrm{e}^{{\bm{x}}_{1}^{*\top}\Gamma_{u}{\bm{x}}/\rho}\right]+\tilde{O}(N_{1}^{-1/2}),$
@@
$\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\mathrm{e}^{y_{i}/\rho}\mathrm{e}^{{\bm{x}}_{i}^{*\top}\Gamma_{u}{\bm{x}}/\rho}=\mathbb{E}_{{\bm{x}}_{1},\zeta_{1}}\left[\mathrm{e}^{\bar{\sigma}_{*}(\left\langle{\bm{x}}_{1},\beta\right\rangle)/\rho+\zeta_{1}/\rho}\mathrm{e}^{{\bm{x}}_{1}^{*\top}\Gamma_{u}{\bm{x}}/\rho}\right]+\tilde{O}(N_{1}^{-1/2}).$
Moreover, noting that $\sqrt{r}\Gamma^{*}{\bm{x}}_{i}^{*}=\tilde{O}(\sqrt{r})$ (see Lemma H.4), from Lemma H.6, we have
$\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\sqrt{r}\Gamma^{*}\frac{y_{i}}{\rho}\mathrm{e}^{y_{i}/\rho}\mathrm{e}^{{\bm{x}}_{i}^{*\top}\Gamma_{u}{\bm{x}}/\rho}{\bm{x}}_{i}^{*}$
$=\sqrt{r}\Gamma^{*}\mathbb{E}_{{\bm{x}}_{1},\zeta_{1}}\left[\frac{\bar{\sigma}_{*}(\left\langle{\bm{x}}_{1},\beta\right\rangle)+\zeta_{1}}{\rho}\mathrm{e}^{\bar{\sigma}_{*}(\left\langle{\bm{x}}_{1},\beta\right\rangle)/\rho+\zeta_{1}/\rho}\mathrm{e}^{{\bm{x}}_{1}^{*\top}\Gamma_{u}{\bm{x}}/\rho}{\bm{x}}_{1}^{*}\right]+\tilde{O}(r^{1/2}N_{1}^{-1/2}),$
@@
$\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\sqrt{r}\Gamma^{*}\mathrm{e}^{y_{i}/\rho}\mathrm{e}^{{\bm{x}}_{i}^{*\top}\Gamma_{u}{\bm{x}}/\rho}{\bm{x}}_{i}^{*}=\sqrt{r}\Gamma^{*}\mathbb{E}_{{\bm{x}}_{1},\zeta_{1}}\left[\mathrm{e}^{\bar{\sigma}_{*}(\left\langle{\bm{x}}_{1},\beta\right\rangle)/\rho+\zeta_{1}/\rho}\mathrm{e}^{{\bm{x}}_{1}^{*\top}\Gamma_{u}{\bm{x}}/\rho}{\bm{x}}_{1}^{*}\right]+\tilde{O}(r^{1/2}N_{1}^{-1/2}).$
@@
$\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\frac{y_{i}}{\rho}\mathrm{e}^{y_{i}/\rho}\mathrm{e}^{{\bm{x}}_{i}^{*\top}\Gamma_{u}{\bm{x}}/\rho}\left\langle{\bm{x}}_{i}^{*},{\bm{u}}\right\rangle$
$=$ $\mathbb{E}_{{\bm{x}}_{1},\zeta_{1}}\left[\frac{\bar{\sigma}_{*}(\left\langle{\bm{x}}_{1},\beta\right\rangle)+\zeta_{1}}{\rho}\mathrm{e}^{\bar{\sigma}_{*}(\left\langle{\bm{x}}_{1},\beta\right\rangle)/\rho+\zeta_{1}/\rho}\mathrm{e}^{{\bm{x}}_{1}^{*\top}\Gamma_{u}{\bm{x}}/\rho}\left\langle{\bm{x}}_{1}^{*},{\bm{u}}\right\rangle\right]+\tilde{O}(C_{u}N_{1}^{-1/2}),$
@@
$\frac{1}{N_{1}}\sum_{i=1}^{N_{1}}\mathrm{e}^{y_{i}/\rho}\mathrm{e}^{{\bm{x}}_{i}^{*\top}\Gamma_{u}{\bm{x}}/\rho}\left\langle{\bm{x}}_{i}^{*},{\bm{u}}\right\rangle=\mathbb{E}_{{\bm{x}}_{1},\zeta_{1}}\left[\mathrm{e}^{\bar{\sigma}_{*}(\left\langle{\bm{x}}_{1},\beta\right\rangle)/\rho+\zeta_{1}/\rho}\mathrm{e}^{{\bm{x}}_{1}^{*\top}\Gamma_{u}{\bm{x}}/\rho}\left\langle{\bm{x}}_{1}^{*},{\bm{u}}\right\rangle\right]+\tilde{O}(C_{u}N_{1}^{-1/2}).$
@@
Moreover, from Lemma C.1, it holds that
@@
where $\sqrt{r}\Gamma^{*}{\bm{x}}=\tilde{O}(\sqrt{r})$ (See Lemma H.4) and $\gamma({\bm{x}},y)=\tilde{O}(1)$ with high probability, which means that $\sqrt{r}\Gamma^{*}\frac{\partial f_{\mathrm{IC}}}{\partial{\bm{u}}}=\tilde{O}_{p}(\alpha mC_{u}\sqrt{r})$. Therefore, we obtain
@@
Moreover, $\mathbb{E}_{{\bm{x}}}[\nu]=0$ by definition, and $\|\nu\|$ has sub-Weibull tail from Lemma C.2. Thus, what remains is to calculate $\mathbb{E}_{{\bm{x}}}[y\cdot\sqrt{r}\Gamma^{*}\nabla_{\bm{u}}{f_{\mathrm{IC}}}({\bm{x}})]$:
@@
and from Lemma H.7, it holds that
@@
with high probability. Finally, we evaluate the term $\mathbb{E}_{{\bm{x}}}[\alpha L_{m}\sigma_{*}(\left\langle\beta,{\bm{x}}\right\rangle)(\sqrt{r}\Gamma^{*}{\bm{x}}\cdot n_{1}+\left\langle{\bm{x}},{\bm{u}}\right\rangle{\bm{n}}_{2})]$. From Corollary 17 of Oko et al. (2024), we have $\sigma_{*}(\left\langle\beta,{\bm{x}}\right\rangle)=O_{p}(\log^{\deg(\sigma_{*})/2}{d})$. Moreover, from Lemma C.1, it holds that $n_{1}=\tilde{O}(C_{u}\sqrt{\frac{1}{N_{1}}})$ and ${\bm{n}}_{2}=\tilde{O}(\sqrt{\frac{r}{N_{1}}})$ with high probability. Therefore, we arrive at